OpenStudy (anonymous):

Suppose that the cost of using solar energy to heat a home involves an initial cost of $10,000 with an average yearly cost of $200 for every year after the first. Suppose also that heating with oil involves an initial cost of $4000 with an average yearly cost of $800 for every year after the first. Use the graph below as a model to graph this information on a separate sheet of paper and answer the questions given. What are the coordinates of the point where the lines intersect?

OpenStudy (anonymous):

Let s(x) represent usage of solar energy: Let o(x) represent usage of oil: s(x)=o(x) where they intersect 10000+200x=4000+800x 6000=600x x=10, find y: 4000+800(10)=10000+200(10)=12000 (10, 12000)

OpenStudy (anonymous):

is it (12,10)

OpenStudy (anonymous):

(10,12000)

OpenStudy (anonymous):

it does't have a answer choice with 1200

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

a) (10,11) b) (11,12) c) (12,11) d) (12,10)

OpenStudy (anonymous):

Are the y-units represented in thousands of dollars?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

OK, i slightly misread the problem. Every year AFTER the first it cost 200x or 800x. So: 10000+200(x-1)=4000+800(x-1) 6000=600(x-1) 6000=600x-600 6600=600x x=11

OpenStudy (anonymous):

I still get 12000 for the y-value. Out of those choices, I'd assume they are representing the y-axis in thousands; I'd go with option B.